If AD=2,DB=x-1,AE=x and EC=x+2, find AE
Found 2 solutions by aliafzal14, JBarnum:
Answer by aliafzal14(13)

If AD=2,DB=x-1,AE=x and EC=x+2, find AE
THIS IS YOUR QUESTION
PLEASE CORRECT YOUR QUESTION BECAUSE THE VALUE WHICH IS USED TO BE FIND HAS A VALUE ALREADY
AE = X

Answer by JBarnum(2146)

Also wanted to add that a triangle only has 3 points and there is nothing in the question to indicate which letter is at a point and if it's triangle abc then where is d and e? So this question is actually completely invalid please make sure you post the question in the correct category and write the question fully and correctly.